OpenStudy (anonymous):

Find the circumference of a circle witha radius of 3 units.. A) 9.42 B) 20.1 units C) 18.84 units D) 15.94 units

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow Circumference = 2\pi r }\)

Parth (parthkohli):

Here, it'd be \(\Large \color{MidnightBlue}{\Rightarrow 2 \times 3.14 \times 3 }\) \(\Large \color{MidnightBlue}{\Rightarrow 6 \times 3.14 }\) \(\Large \color{MidnightBlue}{\Rightarrow 18.84 units}\)

OpenStudy (anonymous):

I believe it's C

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

Oh, yes

OpenStudy (anonymous):

C = Circumference r = radius π = approximate value 3.14 So just plug things in. C = 2(3.14)(3)

Parth (parthkohli):

Exactly